# Isometric immersions into three-dimensional unimodular metric Lie groups

### v1

@@ 4.1. Recovering the immersion from the shape operator
Case A: $\epsilon_{1}c_{1}=\epsilon_{2}c_{2}\neq\epsilon_{3}c_{3}\neq 0$. It comprises all surfaces in $\mathbb{E}(\kappa,\tau)$, in which case we recover Daniel’s fundamental theorem [6], spacelike surfaces in $\mathbb{L}(\kappa,\tau)$, and timelike surfaces in $\widehat{\mathbb{L}}(\kappa,\tau)$. Since the vector field defined by (3.8) reads $T=-c_{3}(\epsilon_{1}c_{1}-\epsilon_{3}c_{3})\nu_{3}T_{3}$ and the last equations in ((((iv)))) and ((((v)))) can be rewritten by taking into account (3.3) and the fact that $\langle X,T_{1}\rangle T_{2}-\langle X,T_{2}\rangle T_{1}=\epsilon_{1}\epsilon_{2}\epsilon_{3}\nu_{3}JX$, we end up with five conditions:
@@ 5. $(\textsc{v}^{*})$
so that we get $\mathrm{d}T_{3}^{\flat}=-c_{3}\nu_{3}\,\omega^{1}\wedge\omega^{2}$. We can then employ this differential and Equations (4.4), (4.5) and (3.12) to compute $\mathrm{d}\Theta$. Also taking into account (3.13), it can be shown that $\mathrm{d}\Theta+\frac{1}{2}[\Theta,\Theta]=0$ after a rather long (but simple) calculation. Here, the Lie bracket is computed in the Lie algebra $\mathfrak{so}^{\epsilon}_{3}(\mathbb{R})$, which is a Lie subalgebra of $\mathfrak{gl}_{3}(\mathbb{R})$, whence $\tfrac{1}{2}[\Theta,\Theta]^{\alpha}_{\beta}=(\Theta\wedge\Theta)^{\alpha}_{\beta}=\sum_{\gamma=1}^{3}\Theta^{\alpha}_{\gamma}\wedge\Theta^{\gamma}_{\beta}$ (i.e., the wedge product has to be computed matrixwise). By means of [30, Thm. 3.7.14], since $\Sigma$ is assumed simply connected, we can ensure that there is smooth map $M:\Sigma\to\mathrm{SO}_{3}^{\epsilon}(\mathbb{R})$ such that $M^{-1}\mathrm{d}M=\Theta=\Omega+L$ and such an $M$ is unique up to left multiplication by a constant matrix $A\in\mathrm{SO}_{3}^{\epsilon}(\mathbb{R})$.
@@
As $\tau\to 0$, the unimodular Lie group structure we have been considering in $\mathbb{E}(\kappa,\tau)$, $\mathbb{L}(\kappa,\tau)$ and $\widehat{\mathbb{L}}(\kappa,\tau)$ disappears. However, these families converge to the product manifolds $\mathbb{M}^{2}(\kappa)\times\mathbb{R}$, $\mathbb{M}^{2}(\kappa)\times\mathbb{R}_{1}$ and $\mathbb{M}^{2}_{1}(\kappa)\times\mathbb{R}$ by just taking limits of the metrics given by (2.11) and (2.12). The fundamental equations (i)-(v) in Theorem 4.2 also have nice limits as $\tau\to 0$ because they ultimately depend on $c_{1}c_{3}=\pm\kappa$ and $c_{3}^{2}=\tau^{2}$. Interestingly, the limit version of Theorem 4.2 still holds true by the work of Daniel in $\mathbb{M}^{n}(\kappa)\times\mathbb{R}$ [7, Thm. 3.3], which was later adapted to the Lorentzian case $\mathbb{M}^{n}(\kappa)\times\mathbb{R}_{1}$ by Roth [29, Thm. 1]. Finally, the case of $\mathbb{M}^{2}_{1}(\kappa)\times\mathbb{R}$ is a particular case of the warped products considered by Lawn–Ortega [15, Thm. 1].
@@ 4.2. Lorentzian versions of the Daniel correspondence
Let $S$ and $\widetilde{S}$ be fields of symmetric operators in a Riemannian or Lorentzian surface $\Sigma$ and assume that $H=\epsilon\operatorname{tr}(S)$ and $\widetilde{H}=\epsilon\operatorname{tr}(\widetilde{S})$ are constants for some $\epsilon\in\{-1,1\}$. If $\widetilde{S}-\epsilon\widetilde{H}\mathrm{id}={\operatorname{Rot}_{\theta}}\circ{(S-\epsilon H\mathrm{id})}$ for some $\theta\in\mathbb{R}$, then the following assertions hold true:
@@ 5. $(\textsc{v}^{*}_{\mathbb{L}})$
Likewise, in the case of $\widehat{\mathbb{L}}(\kappa,\tau)$-spaces, we can set $c_{1}=\frac{-\kappa}{2\tau}$, $c_{2}=\frac{\kappa}{2\tau}$ and $c_{3}=2\tau\neq 0$ with signs $\epsilon_{1}=\epsilon_{3}=1$ and $\epsilon_{3}=1$ in order to get alike equations
  1. $(\textsc{i}^{*}_{\widehat{\mathbb{L}}})$
@@ 5. $(\textsc{v}^{*}_{\widehat{\mathbb{L}}})$
are the fundamental data of another isometric immersion $\widetilde{\phi}:\Sigma\to\widehat{\mathbb{L}}(\widetilde{\kappa},\widetilde{\tau})$ (resp. $\widehat{\mathbb{L}}(\widetilde{\kappa},\widetilde{\tau})$) with cmc $\widetilde{H}$ and the same sign $\hat{\epsilon}_{3}$.
@@ 5. On the generalized Lawson correspondence
It is also interesting to point out the not so well known fact that the Daniel correspondence generalizes the classical Lawson correspondence just partially. When we restrict to $\mathbb{E}(\kappa,\tau)$-spaces with $\kappa-4\tau^{2}=0$, the Daniel correspondence has one degree of freedom (the phase angle determines the target space); on the contrary, the Lawson correspondence has two degrees of freedom. For instance, transforming minimal surfaces in $\mathbb{S}^{3}$ into minimal surfaces in $\mathbb{S}^{3}$ by a rotation of the shape operator is considered within Lawson’s correspondence (see also [3, §5]), but not explained by Daniel’s. It is precisely in those cases in the intersection of Lawson and Daniel in which one can assume that $g$ and $\widetilde{g}$ are globally congruent.
@@
Firstly, Gauss equation ((((i)))) gives $K=\det(S)-\sum_{i=1}^{3}a_{i}\nu_{i}^{2}$, which must be equal to $\det(\widetilde{S})-\sum_{i=1}^{3}\widetilde{a}_{i}\widetilde{\nu}_{i}^{2}$ since $K$ is the intrinsic curvature of $\Sigma$. By item (a) of Lemma 4.4, we get
@@
Secondly, by item (b) of Lemma 4.4 (here we are actually using that $H$ and $\widetilde{H}$ are constants) and (4.8), we can write Codazzi equation ((((ii)))) as
@@
for all $X,Y\in\mathfrak{X}(\Sigma)$, where $T=\sum_{i=1}^{3}a_{i}\nu_{i}T_{i}$ is the vector field given by (3.8). By plugging $X=e_{1}$ and $Y=e_{2}$ into (5.3), where $\{e_{1},e_{2}\}$ is an orthonormal frame of $\Sigma$, we get $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$. In particular, the weaker condition $\|\widetilde{T}\|^{2}=\|T\|^{2}$ yields $\sum_{i,j=1}^{3}a_{i}a_{j}\nu_{i}\nu_{j}\langle T_{i},T_{j}\rangle=\sum_{i,j=1}^{3}\widetilde{a}_{i}\widetilde{a}_{j}\widetilde{\nu}_{i}\widetilde{\nu}_{j}\langle\widetilde{T}_{i},\widetilde{T}_{j}\rangle$, which can be simplified by means of the algebraic relations ((((iii)))) to
@@
On the one hand, it follows from the algebraic relations ((((iii)))) that $\langle\widetilde{T}_{i},\widetilde{T}_{j}\rangle=\langle T_{i},T_{j}\rangle$ for all $i,j\in\{1,2,3\}$, and this means that there is an field of isometric operators $R$ such that $R(T_{i})=\widetilde{T}_{i}$. On the other hand, the vector fields defined by (3.33) only depend on the angle functions and $J$, so they satisfy $\widetilde{X}_{i}=X_{i}$. This also means that the functions defined by (3.37) satisfy $\widetilde{\psi}=\psi$. Since we are assuming that the Gauss map $g$ has rank $2$, Proposition 3.14 tells us that $\psi$ and $\widetilde{\psi}$ are nowhere zero, whence (3.38) implies that $R$ is actually a rotation at each tangent plane of $\Sigma$. (This argument is rather similar to that of the proof item (c) of Theorem 3.8, but now we have to deal with two possibly different ambient spaces.) Since $\sum_{i=1}^{3}\widetilde{\nu}_{i}\widetilde{T}_{i}=0$ and $\widetilde{\nu}_{i}=\nu_{i}$, we get
@@ 5.2. The degenerate case
The last sum clearly equals $-2H$, whereas the other term in the right-hand side of (5.13) swiftly gives $(c_{2}-c_{1})\nu_{1}\nu_{2}$ when expanding ¯$\overline{\nabla}_{e_{i}}E_{3}=\sum_{k=1}^{2}\langle e_{i},E_{k}\rangle\overline{\nabla}_{E_{k}}E_{3}$ and using the ambient connection in (2.4). On the other hand, we can compute
@@
The first sum in the right-hand side of (5.14) vanishes since ¯$\overline{\nabla}_{e_{i}}N$ is proportional to $e_{i}$, whereas the second sum equals $c_{3}\nu_{3}$ by expanding each ¯$\overline{\nabla}_{e_{i}}E_{3}$ as above. ∎
@@ 5.2.1. Case bcv-bcv
Assume first that $(a_{1},a_{3})=(b_{3},b_{1})$. We claim this does not lead to any correspondence. Since $a_{3}-a_{1}=b_{1}-b_{3}$, we infer that $\widetilde{\kappa}-4\widetilde{\tau}^{2}=-(\kappa-4\tau^{2})\neq 0$ ($G$ and $\widetilde{G}$ do not have constant curvature). Consequently, we can rewrite $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ as $\widetilde{\nu}_{3}\widetilde{T}_{3}=-\nu_{3}\operatorname{Rot}_{\theta}(T_{3})$. This subcase gives $\widetilde{\nu}_{3}^{2}=1-\nu_{3}^{2}$, whence $\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3}=-\nu_{3}\nabla\nu_{3}$. Expanding out $\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3}$ by the fundamental equation ((((v)))), we reach
@@ 5.2.2. Case sum-sum
Assume first that $(a_{1},a_{3})=(b_{1},b_{3})$. This swiftly gives $\widetilde{H}^{2}=H^{2}$ and $\widetilde{c}_{1}\widetilde{c}_{2}=c_{1}c_{2}$. We can change $\widetilde{\phi}$ by an ambient isometry to further assume $\widetilde{\nu}_{3}=\nu_{3}$ as in the above case bcv-bcv, which does not alter $\widetilde{T}$. Since $c_{1}c_{2}\neq 0$ (otherwise, we have $\widetilde{c}_{1}\widetilde{c}_{2}=0$ and we can indeed reduce to the case bcv-bcv), the condition $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ amounts to $\widetilde{T}_{3}=\operatorname{Rot}_{\theta}(T_{3})$ at points with $\nu_{3}\neq 0$ by (5.12). By expanding the identity $\langle\nabla\widetilde{\nu}_{3},X\rangle=\langle\nabla\nu_{3},X\rangle$ in the spirit of (5.15) and using ((((v)))), we obtain
@@
  • If $\theta=0$, Equation (5.18) yields $\widetilde{c}_{3}=c_{3}$ and hence $\{\widetilde{c}_{1},\widetilde{c}_{2}\}$ is a permutation of $\{c_{1},c_{2}\}$ (both pairs have the same sum and product), so we can further assume that $\widetilde{c}_{1}=c_{1}$ and $\widetilde{c}_{2}=c_{2}$ by swapping indexes. Multiplying (5.17) by $\widetilde{T}_{3}=T_{3}$ by means of the algebraic relations ((((iii)))), we get
@@
Using the algebraic relations ((((iii)))), we can express
@@ 5.2.3. Case sum-bcv
The subcase $(a_{1},a_{3})=(b_{1},b_{3})$ gives $c_{1}c_{2}+H^{2}=\widetilde{\tau}^{2}+\widetilde{H}^{2}$ and $-c_{1}c_{2}+H^{2}=\widetilde{\kappa}-3\widetilde{\tau}^{2}+\widetilde{H}^{2}$ with $\widetilde{\nu}_{3}=\nu_{3}$ (after a isometry as in previous cases). In particular, we get $\widetilde{\kappa}-4\widetilde{\tau}^{2}=-2c_{1}c_{2}\neq 0$, whence $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ reads $\widetilde{T}_{3}=\operatorname{Rot}_{\theta}(T_{3})$ at points where $\nu_{3}\neq 0$. Taking the divergence by means of Lemma 5.4, we get
@@
Finally, assume $(a_{1},a_{3})=(b_{3},b_{1})$, so that $-c_{1}c_{2}+H^{2}=\widetilde{\tau}^{2}+\widetilde{H}^{2}$ and $c_{1}c_{2}+H^{2}=\widetilde{\kappa}-3\widetilde{\tau}^{2}+\widetilde{H}^{2}$ with $\widetilde{\nu}_{3}^{2}=1-\nu_{3}^{2}$. Since this implies that $\widetilde{\kappa}-4\widetilde{\tau}^{2}=2c_{1}c_{2}\neq 0$, the equality $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ is written as $\widetilde{\nu}_{3}\widetilde{T}_{3}=-\nu_{3}\operatorname{Rot}_{\theta}(T)$ and we also have $\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3}=-\nu_{3}\nabla\nu_{3}$. Proceeding as in the above cases, we can expand $\langle\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3},X\rangle+\langle\nu_{3}\nabla\nu_{3},X\rangle=0$ by means of ((((v)))) to obtain

### v2

@@ 4.1. Recovering the immersion from the shape operator
Case A: $\epsilon_{1}c_{1}=\epsilon_{2}c_{2}\neq\epsilon_{3}c_{3}\neq 0$. It comprises all surfaces in $\mathbb{E}(\kappa,\tau)$, in which case we recover Daniel’s fundamental theorem [7], spacelike surfaces in $\mathbb{L}(\kappa,\tau)$, and timelike surfaces in $\widehat{\mathbb{L}}(\kappa,\tau)$. Since the vector field defined by (3.8) reads $T=-c_{3}(\epsilon_{1}c_{1}-\epsilon_{3}c_{3})\nu_{3}T_{3}$ and the last equations in ((iv)) and ((v)) can be rewritten by taking into account (3.3) and the fact that $\langle X,T_{1}\rangle T_{2}-\langle X,T_{2}\rangle T_{1}=\epsilon_{1}\epsilon_{2}\epsilon_{3}\nu_{3}JX$, we end up with five conditions:
@@ 5. $(\textsc{v}^{*})$
so that we get $\mathrm{d}T_{3}^{\flat}=-c_{3}\nu_{3}\,\omega^{1}\wedge\omega^{2}$. We can then employ this differential and Equations (4.4), (4.5) and (3.12) to compute $\mathrm{d}\Theta$. Also taking into account (3.13), it can be shown that $\mathrm{d}\Theta+\frac{1}{2}[\Theta,\Theta]=0$ after a rather long (but simple) calculation. Here, the Lie bracket is computed in the Lie algebra $\mathfrak{so}^{\epsilon}_{3}(\mathbb{R})$, which is a Lie subalgebra of $\mathfrak{gl}_{3}(\mathbb{R})$, whence $\tfrac{1}{2}[\Theta,\Theta]^{\alpha}_{\beta}=(\Theta\wedge\Theta)^{\alpha}_{\beta}=\sum_{\gamma=1}^{3}\Theta^{\alpha}_{\gamma}\wedge\Theta^{\gamma}_{\beta}$ (i.e., the wedge product has to be computed matrixwise). By means of [31, Thm. 3.7.14], since $\Sigma$ is assumed simply connected, we can ensure that there is smooth map $M:\Sigma\to\mathrm{SO}_{3}^{\epsilon}(\mathbb{R})$ such that $M^{-1}\mathrm{d}M=\Theta=\Omega+L$ and such an $M$ is unique up to left multiplication by a constant matrix $A\in\mathrm{SO}_{3}^{\epsilon}(\mathbb{R})$.
@@
As $\tau\to 0$, the unimodular Lie group structure we have been considering in $\mathbb{E}(\kappa,\tau)$, $\mathbb{L}(\kappa,\tau)$ and $\widehat{\mathbb{L}}(\kappa,\tau)$ disappears. However, these families converge to the product manifolds $\mathbb{M}^{2}(\kappa)\times\mathbb{R}$, $\mathbb{M}^{2}(\kappa)\times\mathbb{R}_{1}$ and $\mathbb{M}^{2}_{1}(\kappa)\times\mathbb{R}$ by just taking limits of the metrics given by (2.11) and (2.12). The fundamental equations (i)-(v) in Theorem 4.2 also have nice limits as $\tau\to 0$ because they ultimately depend on $c_{1}c_{3}=\pm\kappa$ and $c_{3}^{2}=\tau^{2}$. Interestingly, the limit version of Theorem 4.2 still holds true by the work of Daniel in $\mathbb{M}^{n}(\kappa)\times\mathbb{R}$ [8, Thm. 3.3], which was later adapted to the Lorentzian case $\mathbb{M}^{n}(\kappa)\times\mathbb{R}_{1}$ by Roth [30, Thm. 1]. Finally, the case of $\mathbb{M}^{2}_{1}(\kappa)\times\mathbb{R}$ is a particular case of the warped products considered by Lawn–Ortega [16, Thm. 1].
@@ 4.2. Lorentzian versions of the Daniel correspondence
[…]
@@ 5. $(\textsc{v}^{*}_{\mathbb{L}})$
Likewise, in the case of $\widehat{\mathbb{L}}(\kappa,\tau)$-spaces, we can set $c_{1}=\frac{-\kappa}{2\tau}$, $c_{2}=\frac{\kappa}{2\tau}$ and $c_{3}=2\tau\neq 0$ with signs $\epsilon_{1}=\epsilon_{3}=1$ and $\epsilon_{2}=-1$ in order to get alike equations
[…]
@@
[…]
@@ 5. On the generalized Lawson correspondence
It is also interesting to point out the not so well known fact that the Daniel correspondence generalizes the classical Lawson correspondence just partially. When we restrict to $\mathbb{E}(\kappa,\tau)$-spaces with $\kappa-4\tau^{2}=0$, the Daniel correspondence has one degree of freedom (the phase angle determines the target space); on the contrary, the Lawson correspondence has two degrees of freedom. For instance, transforming minimal surfaces in $\mathbb{S}^{3}$ into minimal surfaces in $\mathbb{S}^{3}$ by a rotation of the shape operator is considered within Lawson’s correspondence (see also [5, §5]), but not explained by Daniel’s. It is precisely in those cases in the intersection of Lawson and Daniel in which one can assume that $g$ and $\widetilde{g}$ are globally congruent.
@@
Firstly, Gauss equation ((i)) gives $K=\det(S)-\sum_{i=1}^{3}a_{i}\nu_{i}^{2}$, which must be equal to $\det(\widetilde{S})-\sum_{i=1}^{3}\widetilde{a}_{i}\widetilde{\nu}_{i}^{2}$ since $K$ is the intrinsic curvature of $\Sigma$. By item (a) of Lemma 4.4, we get
@@
Secondly, by item (b) of Lemma 4.4 (here we are actually using that $H$ and $\widetilde{H}$ are constants) and (4.8), we can write Codazzi equation ((ii)) as
@@
for all $X,Y\in\mathfrak{X}(\Sigma)$, where $T=\sum_{i=1}^{3}a_{i}\nu_{i}T_{i}$ is the vector field given by (3.8). By plugging $X=e_{1}$ and $Y=e_{2}$ into (5.3), where $\{e_{1},e_{2}\}$ is an orthonormal frame of $\Sigma$, we get $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$. In particular, the weaker condition $\|\widetilde{T}\|^{2}=\|T\|^{2}$ yields $\sum_{i,j=1}^{3}a_{i}a_{j}\nu_{i}\nu_{j}\langle T_{i},T_{j}\rangle=\sum_{i,j=1}^{3}\widetilde{a}_{i}\widetilde{a}_{j}\widetilde{\nu}_{i}\widetilde{\nu}_{j}\langle\widetilde{T}_{i},\widetilde{T}_{j}\rangle$, which can be simplified by means of the algebraic relations ((iii)) to
@@
On the one hand, it follows from the algebraic relations ((iii)) that $\langle\widetilde{T}_{i},\widetilde{T}_{j}\rangle=\langle T_{i},T_{j}\rangle$ for all $i,j\in\{1,2,3\}$, and this means that there is an field of isometric operators $R$ such that $R(T_{i})=\widetilde{T}_{i}$. On the other hand, the vector fields defined by (3.33) only depend on the angle functions and $J$, so they satisfy $\widetilde{X}_{i}=X_{i}$. This also means that the functions defined by (3.37) satisfy $\widetilde{\psi}=\psi$. Since we are assuming that the Gauss map $g$ has rank $2$, Proposition 3.14 tells us that $\psi$ and $\widetilde{\psi}$ are nowhere zero, whence (3.38) implies that $R$ is actually a rotation at each tangent plane of $\Sigma$. (This argument is rather similar to that of the proof item (c) of Theorem 3.8, but now we have to deal with two possibly different ambient spaces.) Since $\sum_{i=1}^{3}\widetilde{\nu}_{i}\widetilde{T}_{i}=0$ and $\widetilde{\nu}_{i}=\nu_{i}$, we get
@@ 5.2. The degenerate case
The last sum clearly equals 2H-2H, whereas the other term in the right-hand side of (5.13) swiftly gives $(c_{2}-c_{1})\nu_{1}\nu_{2}$ when expanding $\overline{\nabla}_{e_{i}}E_{3}=\sum_{k=1}^{2}\langle e_{i},E_{k}\rangle\overline{\nabla}_{E_{k}}E_{3}$ and using the ambient connection in (2.4). On the other hand, we can compute
@@
The first sum in the right-hand side of (5.14) vanishes since $\overline{\nabla}_{e_{i}}N$ is proportional to $e_{i}$, whereas the second sum equals $c_{3}\nu_{3}$ by expanding each $\overline{\nabla}_{e_{i}}E_{3}$ as above. ∎
@@ 5.2.1. Case bcv-bcv
Assume first that $(a_{1},a_{3})=(b_{3},b_{1})$. We claim this does not lead to any correspondence. Since $a_{3}-a_{1}=b_{1}-b_{3}$, we infer that $\widetilde{\kappa}-4\widetilde{\tau}^{2}=-(\kappa-4\tau^{2})\neq 0$ ($G$ and $\widetilde{G}$ do not have constant curvature). Consequently, we can rewrite $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ as $\widetilde{\nu}_{3}\widetilde{T}_{3}=-\nu_{3}\operatorname{Rot}_{\theta}(T_{3})$. This subcase gives $\widetilde{\nu}_{3}^{2}=1-\nu_{3}^{2}$, whence $\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3}=-\nu_{3}\nabla\nu_{3}$. Expanding out $\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3}$ by the fundamental equation ((v)), we reach
@@ 5.2.2. Case sum-sum
Assume first that $(a_{1},a_{3})=(b_{1},b_{3})$. This swiftly gives $\widetilde{H}^{2}=H^{2}$ and $\widetilde{c}_{1}\widetilde{c}_{2}=c_{1}c_{2}$. We can change $\widetilde{\phi}$ by an ambient isometry to further assume $\widetilde{\nu}_{3}=\nu_{3}$ as in the above case bcv-bcv, which does not alter $\widetilde{T}$. Since $c_{1}c_{2}\neq 0$ (otherwise, we have $\widetilde{c}_{1}\widetilde{c}_{2}=0$ and we can indeed reduce to the case bcv-bcv), the condition $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ amounts to $\widetilde{T}_{3}=\operatorname{Rot}_{\theta}(T_{3})$ at points with $\nu_{3}\neq 0$ by (5.12). By expanding the identity $\langle\nabla\widetilde{\nu}_{3},X\rangle=\langle\nabla\nu_{3},X\rangle$ in the spirit of (5.15) and using ((v)), we obtain
@@
  • If $\theta=0$, Equation (5.18) yields $\widetilde{c}_{3}=c_{3}$ and hence $\{\widetilde{c}_{1},\widetilde{c}_{2}\}$ is a permutation of $\{c_{1},c_{2}\}$ (both pairs have the same sum and product), so we can further assume that $\widetilde{c}_{1}=c_{1}$ and $\widetilde{c}_{2}=c_{2}$ by swapping indexes. Multiplying (5.17) by $\widetilde{T}_{3}=T_{3}$ by means of the algebraic relations ((iii)), we get
@@
Using the algebraic relations ((iii)), we can express
@@ 5.2.3. Case sum-bcv
The subcase $(a_{1},a_{3})=(b_{1},b_{3})$ gives $c_{1}c_{2}+H^{2}=\widetilde{\tau}^{2}+\widetilde{H}^{2}$ and $c_{1}c_{2}+H^{2}=\widetilde{\kappa}-3\widetilde{\tau}^{2}+\widetilde{H}^{2}$ with $\widetilde{\nu}_{3}=\nu_{3}$ (after a isometry as in previous cases). In particular, we get $\widetilde{\kappa}-4\widetilde{\tau}^{2}=-2c_{1}c_{2}\neq 0$, whence $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ reads $\widetilde{T}_{3}=\operatorname{Rot}_{\theta}(T_{3})$ at points where $\nu_{3}\neq 0$. Taking the divergence by means of Lemma 5.4, we get
@@
Finally, assume $(a_{1},a_{3})=(b_{3},b_{1})$, so that $c_{1}c_{2}+H^{2}=\widetilde{\tau}^{2}+\widetilde{H}^{2}$ and $c_{1}c_{2}+H^{2}=\widetilde{\kappa}-3\widetilde{\tau}^{2}+\widetilde{H}^{2}$ with $\widetilde{\nu}_{3}^{2}=1-\nu_{3}^{2}$. Since this implies that $\widetilde{\kappa}-4\widetilde{\tau}^{2}=2c_{1}c_{2}\neq 0$, the equality $\widetilde{T}=\operatorname{Rot}_{\theta}(T)$ is written as $\widetilde{\nu}_{3}\widetilde{T}_{3}=-\nu_{3}\operatorname{Rot}_{\theta}(T)$ and we also have $\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3}=-\nu_{3}\nabla\nu_{3}$. Proceeding as in the above cases, we can expand $\langle\widetilde{\nu}_{3}\nabla\widetilde{\nu}_{3},X\rangle+\langle\nu_{3}\nabla\nu_{3},X\rangle=0$ by means of ((v)) to obtain